OpenStudy (psycho45):

resoudre pour x 4/5(15x+20)−7x=5/6(12x−24)+6

OpenStudy (psycho45):

\[\frac{ 4 }{ 5 }(15x+20)-7x=\frac{ 5 }{ 6 }(12x-24)+6\]

OpenStudy (mathmate):

Distribue: \(12x+16-7x=10x-20+6\) Isole x et simplifie afin de résoudre pour x.

OpenStudy (psycho45):

Merci

OpenStudy (mathmate):

Je vous en prie! :)

OpenStudy (psycho45):

Aïe qui fait mal